Session: May 1998 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (roma at Sat May 30 11:00:18 METDST 1998) Minor failures
sorry lost first 2 scan, not other problems, sky very clody at start and clear at stop.

Torun (K. Borkowski at Sat May 30 12:39:12 METDST 1998) Success
No known problems.

Westerbork (Tony Foley at Sat May 30 14:01:22 METDST 1998) Minor failures
Data looks OK, but still problems with decoder.

Onsala (A. Polatidis at Sun May 31 14:03:44 METDST 1998) Success
No problems with the experiment. Summer weather at last.

Jodrell (Alastair Gunn at Mon Jun 1 15:41:21 METDST 1998) Success
No Known problems.

Effelsberg (dgraham at Tue Jun 2 18:33:09 METDST 1998) Minor failures

Calibration (Jean\-Francois Desmurs at Mon Jun 8 13:24:03 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name v030w.antab. Read the comments in the file to get further information.

Noto (C.Trigilio at Mon Jun 8 17:17:10 METDST 1998) Minor failures
Some tracking problem caused the telescope going on and off very often around 02:53 and around 07:48